OpenStudy (anonymous):

pam is playing with red and black marbles. The number of red marbles she has is 3 more than twice the number of black marbles. She has 42 marbles in all. How many red marbles does she have? A) 13 B)15 C) 29 D) 33

OpenStudy (kropot72):

Let r = number of red Let b = number of black r + b = 42 .........1) b = 42 - r ..........2) 2b + 3 = r ..........3) Substituting the value of b found in 2) into 3) 2(42 - r) + 3 = r 84 - 2r + 3 = r 3r = 87 Can you find the value of r now?

OpenStudy (anonymous):

29

OpenStudy (kropot72):

That's right! Hope you understand how to do it :)